# StreamingThinker: Large Language Models Can Think While Reading

### v1

@@ 1 Introduction
Large language models (LLMs) have shown impressive reasoning capabilities, as exemplified by systems like OpenAI-o1 (Jaech et al., 2024) and DeepSeek-R1 (Guo et al., 2025). Yet, most current approaches follow a batch thinking paradigm, in which reasoning begins only after the entire input context has been received. This paradigm is problematic in scenarios that demand timely responses or dynamic information processing. First, waiting for the full input before reasoning introduces unnecessary latency. Second, as the input increases, attention to earlier information becomes diluted due to the growing disconnect between reasoning steps and their relevant context (Liu et al., 2024; Levy et al., 2024; Zhang et al., 2025b). This weakens coherence and raises the risk of hallucination. To compensate, LLMs often rely on longer chains of thought (CoT) (Wei et al., 2022; Yeo et al., 2025; Chen et al., 2025b) or repeated self-refinement (Wang et al., 2023; Ling et al., 2023; Madaan et al., 2023) to re-focus, which in turn raise computational costs and inflate token usage.
In contrast, human reasoning often unfolds in an immediate and streaming manner. Research in psychology and cognitive science shows that during reading, humans process incoming information instantaneously, including text decoding, meaning construction, background knowledge activation, integrative reasoning, and actively generating inferences to build a coherent understanding (Kintsch, 1988; Graesser et al., 1994). This “thinking while reading” mechanism not only enhances processing efficiency, but also allows reasoning to occur closely alongside the relevant context, minimizing cognitive lag and mitigating the risk of coherence degradation.
@@
Accordingly, we propose StreamingThinker, a framework that instantiates the streaming thinking paradigm. The StreamingThinker integrates a streaming CoT generation pipeline with training and inference frameworks that adapt LLMs to the streaming paradigm. The generation pipeline inserts boundary tokens to the inputs to define minimal reasoning units, prompting LLMs to generate serialized reasoning segments for each unit that are reconstructed into incremental content, filtered by quality evaluation, with reasoning depth controlled through token intervention (Wu et al., 2025). To support streaming training, StreamingThinker introduces two modifications: a streaming attention mask that restricts each reasoning step to past and current input, and a streaming position encoding that independently indexes input and reasoning tokens from zero to eliminate positional contention, ensuring alignment with associated inputs. For inference, StreamingThinker employs parallel KV caches that decouple input encoding from reasoning generation and merge only during cross-attention, enabling true thinking while reading. We conduct a comprehensive evaluation on Qwen3 model family (Yang et al., 2025), covering diverse tasks such as math reasoning, logical reasoning, and context-based QA reasoning. Experimental results indicate that StreamingThinker achieves reasoning performance on par with batch thinking, yet reduces token-level waiting before reasoning by 80% and overall answer latency by over 60%.
@@ 2 Streaming Thinking Paradigm
Human streaming cognition involves two complementary processes: rapidly generating and updating representations as input arrives, and subsequently performing global integration to transform local, shallow understanding into holistic, deep comprehension (Kintsch, 1988).
@@
In the batch thinking paradigm, all input is available simultaneously, so the order of context and question is often overlooked. Yet prior work shows that their relative order can influence reasoning (Chen et al., 2024b; Wei et al., 2024; Xie, 2024), an effect amplified in streaming scenarios. In human reading, two natural input orders commonly occur. In the first, the question is presented before the context, enabling the reader to establish targeted associations as subsequent information is processed. In the second, the context precedes the question, in which case the question remains unavailable during streaming reasoning and the reader must rely solely on the context itself to construct plausible inferences. To approximate these scenarios, our streaming thinking paradigm explicitly distinguishes between the two orders and examines their impact on model reasoning. Some examples are provided in Appendix B.
@@ 3.1 Streaming CoT Generation
To further enforce sequential alignment, a higher-parameter teacher model reconstructs the generated reasoning. Once all sentence-level reasoning traces are generated, they are evaluated using granularity score and sequential consistency score. Passing samples are enhanced with token-level intervention to generate depth-controlled reasoning variants. Samples that fail the evaluation are regenerated, and those still failing under Pass@2 (Chen et al., 2021) metric are discarded.
@@
We propose two evaluation metrics: the granularity score measures the fine-grained alignment of the streaming CoT, while the sequential consistency score assesses whether the reasoning proceeds in a streaming, order-preserving way. The granularity score is defined as the ratio between the number of boundary tokens in the input and those in the output: , where and denote the counts of boundary tokens in the input and output, respectively. When the granularity score is equal to 1, the reasoning matches the input in boundary count, suggesting ideal alignment. The consistency score is defined as the similarity between the input sentence and reasoning sentences, i.e., , where and denote the embedding vectors of the reasoning sentences and the input sentence , respectively. We use SentenceBERT (Reimers & Gurevych, 2019) for sequential consistency calculation.333In cases where a single input sentence corresponds to multiple reasoning sentences, we regard them collectively as one segment. We provide the similarity map between the input and the reasoning in the Appendix C.5.
@@ 3.2 Streaming Training Framework
The RoPE (Su et al., 2024) of LLMs represents relative positions by rotating queries and keys, with the attention between a reasoning token and an input token expressed as , where is the input length and are their positional IDs, and is the rotary matrix. However, in streaming scenarios, the concurrent generation of output and reception of input induces positional contention in the encoding space (Tong et al., 2025; Guo et al., 2024). To address this issue, we assign independent position IDs to input and reasoning tokens, both starting from zero. Formally, the positional IDs of reasoning token and input token are both set to , yielding . This design removes positional contention in streaming parallel processing. Furthermore, identical position IDs ensure that, during streaming reasoning, a reasoning sentence is positioned nearest to its associated input sentence and distant from others, which conforms to the essential principle of streaming alignment.
@@ 4.1 Experimental Settings
To evaluate StreamingThinker, we conduct a comprehensive assessment across three representative reasoning tasks: math reasoning, logical reasoning, and context-based QA reasoning. For math reasoning, we selected GSM-symbolic (Mirzadeh et al., 2025) and MetamathQA (Yu et al., 2024). For logical reasoning, we utilized LogicNLI (Tian et al., 2021) and ProofWriter (Tafjord et al., 2020). Finally, for context-based QA reasoning, PubMedQA (Jin et al., 2019) and HotpotQA (Yang et al., 2018) were employed. All datasets were partitioned into dedicated training and testing sets, with detailed specifications provided in the Appendix D.
@@
The evaluation of streaming scenarios can be viewed as a trade-off between performance and latency. For reasoning performance, we adopt Pass@1 score as the accuracy metric to measure the model’s ability to successfully solve problems. Latency is assessed at two levels: token latency and time latency. At the token level, we use token-to-first-token (TTFT) to measure how many input tokens must be observed before the model begins reasoning.444This is similar with time-to-first-token, but measured in terms of token count. For time latency, we set the LLM’s input speed as the average human speaking rate about 150 words/s (Geva & Yaghoub Zadeh, 2006; Jacewicz et al., 2010), and define the waiting time until the first answer token as the latency.555Detailed definitions and additional evaluation metrics are provided in Appendix F.
@@ 4.2 Effectiveness of the Streaming Thinking Paradigm for LLMs
The attention visualizations in Figures 4(c) and (d) provide additional insights. While the original RoPE exhibits no clear positional preference across the input context, Streaming RoPE shows a pronounced diagonal concentration, reflecting stronger focus on the current context. This bias aligns well with the motivation of streaming thinking: reasoning should primarily rely on information already observed, thereby enabling the model to “think while reading”.
@@ 4.4 Ordering of Context and Question for StreamingThinker
Figure 5 highlights the differences between the two settings. When the question appears first, the model is aware of the reasoning target and tends to expand its logic once relevant context is observed. This leads to higher token usage at depth D1, and since part of the reasoning is already completed during the streaming process, the incremental token growth at deeper levels (D2 and D3) becomes smaller compared to the context-first setting. In contrast, when the context appears first, the model lacks knowledge of which information is salient. As a result, its reasoning remains more conservative, proceeding sentence by sentence without extensive elaboration, which produces lower token usage at D1. Once the question becomes available, later reasoning stages can selectively refine the analysis, bringing accuracy gains by increases in token cost.777The model’s conservative reasoning stems from our cautious data generation strategy—when the question is unknown, excessive logical expansion may cause overthinking.
@@
| Batch-thinking | 0.483 | 1487.56 | 350.08 | 48.35 | 0.566 | 1492.41 | 1485.55 | 48.50 | 0.593 | 1320.36 | 357.47 | 42.90 |
| Streaming, D1 | 0.476 | 604.84 | 5.60 | 0.89 | 0.532 | 318.42 | 46.39 | 1.62 | 0.562 | 298.51 | 31.49 | 1.12 |
| Streaming, D2 | 0.613 | 927.35 | 5.60 | 10.24 | 0.575 | 502.76 | 46.39 | 5.90 | 0.628 | 475.63 | 31.49 | 5.54 |
| Streaming, D3 | 0.625 | 1122.32 | 5.60 | 16.55 | 0.591 | 694.82 | 46.39 | 12.14 | 0.640 | 687.45 | 31.49 | 12.43 |
@@ 5 Related Work
Prior studies on efficient reasoning in LLMs have mainly focused on three directions: token compression (Aggarwal & Welleck, 2025; Xia et al., 2025; Zhang et al., 2025a), structural quantization and pruning (Liu et al., 2025; Srivastava et al., 2025; Zhao et al., 2025), and efficient decoding (Pan et al., 2025; Liao et al., 2025). Token compression methods such as Tokenskip (Xia et al., 2025) condense CoT into fewer tokens to reduce the inference cost. Structural approaches leverage quantization or pruning to compress model parameters for efficient reasoning such as (Srivastava et al., 2025). Efficient decoding has been explored through parallel sampling of generation paths (Pan et al., 2025) or by using smaller models for speculative prediction (Liao et al., 2025) to reduce inference latency. In contrast, our work introduces streaming processing as a complementary dimension of efficiency, allowing reasoning to proceed concurrently with input processing to reduce response latency.
@@
Recent research on streaming LLMs has focused on three directions: architecture adaptation (Tong et al., 2025; Raffel et al., 2024; Guo et al., 2024), latency control (Ma et al., 2018; Ahmed et al., 2025; Cheng et al., 2025), and modality adaptation (Chen et al., 2024a; 2025a; Xie & Wu, 2024). Architecture adaptation addresses the mismatch between decoder-only transformers and streaming settings (Tong et al., 2025), where issues such as positional interference and redundant re-encoding are mitigated through recurrent states, and modified attention mechanisms. Latency control emphasizes fine-grained alignment between input and output, ranging from fixed policies (e.g., wait-k (Ma et al., 2018)) to adaptive scheduling that optimize the accuracy–latency balance. Modality adaptation extends streaming LLMs beyond text to speech recognition (Guo et al., 2024), speech translation (Cheng et al., 2025), and video understanding (Chen et al., 2024a) by integrating modality-specific encoders and synchronization mechanisms. From a reasoning perspective, recent studies (Xie et al., 2025a; Chiang et al., 2025; Xie et al., 2025b) have explored alternating reasoning and generation to approximate streaming operation. Our work differs by explicitly modeling thinking while reading, enabling reasoning to evolve concurrently with incremental input.
@@ 6 Conclusion
Appendix A Prospective Applications of Streaming Thinking
@@
The streaming thinking paradigm is also particularly well-suited for the challenges presented by embodied AI. An embodied agent, like a robot, operates within a dynamic physical environment, which distinguishes it from models that process static information. It continually receives multi-modal sensory data (from cameras, LiDAR, tactile sensors) and is required to respond to this information in a timely manner. In this context, streaming thinking allows the agent to engage in continuous perceptual reasoning, interpreting incoming data to consistently update its internal model of the world. This capability supports dynamic instrumental reasoning, where the model can flexibly plan and re-plan its actions to navigate changing conditions and work towards a goal. For example, a household robot would need to reason about multiple factors at once—such as the movement of a person, the delicacy of an object it plans to handle, and the best path through a cluttered space—while adapting its motor commands accordingly. This close coupling of perception, reasoning, and action in a real-time loop is a core aspect of embodied cognition, and achieving it presents significant challenges for a conventional batch thinking approach.

### v2

@@ 1 Introduction
Large language models (LLMs) have shown impressive reasoning capabilities, as exemplified by systems like OpenAI-o1 (jaech2024openai) and DeepSeek-R1 (guo2025deepseek). Yet, most current approaches follow a batch thinking paradigm, in which reasoning begins only after the entire input context has been received. This paradigm is problematic in scenarios that demand timely responses or dynamic information processing. First, waiting for the full input before reasoning introduces unnecessary latency. Second, as the input increases, attention to earlier information becomes diluted due to the growing disconnect between reasoning steps and their relevant context (liu2024lost; levy2024same; zhang2025attention). This weakens coherence and raises the risk of hallucination. To compensate, LLMs often rely on longer chains of thought (CoT) (wei2022chain; yeo2025demystifying; chen2025towards) or repeated self-refinement (wangself; ling2023deductive; madaan2023self) to re-focus, which in turn raise computational costs and inflate token usage.
In contrast, human reasoning often unfolds in an immediate and streaming manner. Research in psychology and cognitive science shows that during reading, humans process incoming information instantaneously, including text decoding, meaning construction, background knowledge activation, integrative reasoning, and actively generating inferences to build a coherent understanding (kintsch1988role; graesser1994constructing). This “thinking while reading” mechanism not only enhances processing efficiency, but also allows reasoning to occur closely alongside the relevant context, minimizing cognitive lag and mitigating the risk of coherence degradation.
@@
Accordingly, we propose StreamingThinker, a framework that instantiates the streaming thinking paradigm. The StreamingThinker integrates a streaming CoT generation pipeline with training and inference frameworks that adapt LLMs to the streaming paradigm. The generation pipeline inserts boundary tokens to the inputs to define minimal reasoning units, prompting LLMs to generate serialized reasoning segments for each unit that are reconstructed into incremental content, filtered by quality evaluation, with reasoning depth controlled through token intervention (wu2025effectively). To support streaming training, StreamingThinker introduces two modifications: a streaming attention mask that restricts each reasoning step to past and current input, and a streaming position encoding that independently indexes input and reasoning tokens from zero to eliminate positional contention, ensuring alignment with associated inputs. For inference, StreamingThinker employs parallel KV caches that decouple input encoding from reasoning generation and merge only during cross-attention, enabling true thinking while reading. We conduct a comprehensive evaluation on Qwen3 model family (yang2025qwen3), covering diverse tasks such as math reasoning, logical reasoning, and context-based QA reasoning. Experimental results indicate that StreamingThinker achieves reasoning performance on par with batch thinking, yet reduces token-level waiting before reasoning by 80% and overall answer latency by over 60%.
@@ 2 Streaming Thinking Paradigm
Human streaming cognition involves two complementary processes: rapidly generating and updating representations as input arrives, and subsequently performing global integration to transform local, shallow understanding into holistic, deep comprehension (kintsch1988role).
@@
In the batch thinking paradigm, all input is available simultaneously, so the order of context and question is often overlooked. Yet prior work shows that their relative order can influence reasoning (chen2024premise; wei2024unveiling; xie2024order), an effect amplified in streaming scenarios. In human reading, two natural input orders commonly occur. In the first, the question is presented before the context, enabling the reader to establish targeted associations as subsequent information is processed. In the second, the context precedes the question, in which case the question remains unavailable during streaming reasoning and the reader must rely solely on the context itself to construct plausible inferences. To approximate these scenarios, our streaming thinking paradigm explicitly distinguishes between the two orders and examines their impact on model reasoning. Some examples are provided in Appendix B.
@@ 3.1 Streaming CoT Generation
To further enforce sequential alignment, a higher-parameter teacher model reconstructs the generated reasoning. Once all sentence-level reasoning traces are generated, they are evaluated using granularity score and sequential consistency score. Passing samples are enhanced with token-level intervention to generate depth-controlled reasoning variants. Samples that fail the evaluation are regenerated, and those still failing under Pass@2 (chen2021evaluating) metric are discarded.
@@
We propose two evaluation metrics: the granularity score measures the fine-grained alignment of the streaming CoT, while the sequential consistency score assesses whether the reasoning proceeds in a streaming, order-preserving way. The granularity score is defined as the ratio between the number of boundary tokens in the input and those in the output: , where and denote the counts of boundary tokens in the input and output, respectively. When the granularity score is equal to 1, the reasoning matches the input in boundary count, suggesting ideal alignment. The consistency score is defined as the similarity between the input sentence and reasoning sentences, i.e., , where and denote the embedding vectors of the reasoning sentences and the input sentence , respectively. We use SentenceBERT (reimers2019sentence) for sequential consistency calculation.333In cases where a single input sentence corresponds to multiple reasoning sentences, we regard them collectively as one segment. We provide the similarity map between the input and the reasoning in the Appendix C.5.
@@ 3.2 Streaming Training Framework
The RoPE (su2024roformer) of LLMs represents relative positions by rotating queries and keys, with the attention between a reasoning token and an input token expressed as , where is the input length and are their positional IDs, and is the rotary matrix. However, in streaming scenarios, the concurrent generation of output and reception of input induces positional contention in the encoding space (tong2025llm; guo2024decoder). To address this issue, we assign independent position IDs to input and reasoning tokens, both starting from zero. Formally, the positional IDs of reasoning token and input token are both set to , yielding . This design removes positional contention in streaming parallel processing. Furthermore, identical position IDs ensure that, during streaming reasoning, a reasoning sentence is positioned nearest to its associated input sentence and distant from others, which conforms to the essential principle of streaming alignment.
@@ 4.1 Experimental Settings
To evaluate StreamingThinker, we conduct a comprehensive assessment across three representative reasoning tasks: math reasoning, logical reasoning, and context-based QA reasoning. For math reasoning, we selected GSM-symbolic (mirzadehgsm) and MetamathQA (yumetamath). For logical reasoning, we utilized LogicNLI (tian2021diagnosing) and ProofWriter (tafjord2020proofwriter). Finally, for context-based QA reasoning, PubMedQA (jin2019pubmedqa) and HotpotQA (yang2018hotpotqa) were employed. All datasets were partitioned into dedicated training and testing sets, with detailed specifications provided in the Appendix D.
@@
The evaluation of streaming scenarios can be viewed as a trade-off between performance and latency. For reasoning performance, we adopt Pass@1 score as the accuracy metric to measure the model’s ability to successfully solve problems. Latency is assessed at two levels: token latency and time latency. At the token level, we use token-to-first-token (TTFT) to measure how many input tokens must be observed before the model begins reasoning.444This is similar with time-to-first-token, but measured in terms of token count. For time latency, we set the LLM’s input speed as the average human speaking rate about 150 words per minute (geva2006reading; jacewicz2010between), and define the waiting time until the first answer token as the latency.555Detailed definitions and additional evaluation metrics are provided in Appendix F.
@@ 4.2 Effectiveness of the Streaming Thinking Paradigm for LLMs
The attention visualizations in Figures 4(c) and (d) provide additional insights. While the original RoPE exhibits no clear positional preference across the input context, Streaming RoPE shows a pronounced diagonal concentration, reflecting stronger focus on the current context. This bias aligns well with the motivation of streaming thinking: reasoning should primarily rely on information already observed, thereby enabling the model to “think while reading.”
@@ 4.4 Ordering of Context and Question for StreamingThinker
Figure 5 highlights the differences between the two settings.777For clarity, we report the average performance for tasks of the same type. When the question appears first, the model is aware of the reasoning target. This is particularly advantageous in Context-based QA tasks where critical information is sparse; the prior knowledge of the question allows the model to precisely capture key evidence, thereby avoiding the generation of reasoning for irrelevant context (in contrast to domains with denser information). However, for the identified relevant segments, the model tends to expand its logic immediately. This leads to higher token usage at depth D1, and since part of the reasoning is already completed during the streaming process, the incremental token growth at deeper levels (D2 and D3) becomes smaller compared to the context-first setting. In contrast, when the context appears first, the model lacks knowledge of which information is salient. As a result, its reasoning remains more conservative, proceeding sentence by sentence without extensive elaboration, which produces lower token usage at D1 when processing the same context sentence.888The model’s conservative reasoning stems from our cautious data generation strategy—when the question is unknown, excessive logical expansion may cause overthinking. However, this characteristic leads to inefficiency in sparse scenarios (such as the context-based QA tasks): due to the conservative strategy, the model fails to skip irrelevant information, resulting in a significantly longer total generation length.
@@
| Batch, Original | 0.483 | 1487.56 | 350.08 | 48.35 | 0.566 | 1492.41 | 1485.55 | 48.50 | 0.593 | 1320.36 | 357.47 | 42.90 |
| Streaming, D1 | 0.476 | 604.84 | 5.60 | 0.89 | 0.532 | 1172.36 | 46.39 | 1.62 | 0.562 | 298.51 | 31.49 | 1.12 |
| Streaming, D2 | 0.613 | 927.35 | 5.60 | 10.24 | 0.575 | 1266.41 | 46.39 | 5.90 | 0.628 | 475.63 | 31.49 | 5.54 |
| Streaming, D3 | 0.625 | 1122.32 | 5.60 | 16.55 | 0.591 | 1381.73 | 46.39 | 12.14 | 0.640 | 687.45 | 31.49 | 12.43 |
5 Discussion
Efficiency Analysis
We evaluate the model efficiency on Qwen3-4B using 100 samples from GSM-Symbolic dataset. As shown in Table 4, the Streaming paradigm reduces the first-token latency (measured by time consumption) from 28.00s to 6.23s (4.5 speedup). Crucially, the parallel KV cache operations introduce negligible temporal overhead, with and taking less than 5ms combined. Additionally, peak memory usage remains consistent with the baseline (7.99 GB). While bandwidth cost increases, this is primarily due to the inherent multiple prefill phases in streaming scenarios rather than the parallel KV cache mechanism itself. Note that Table 4 reports the latency for each stage separately; the actual end-to-end wall-clock time, which benefits from the concurrent execution of reading and reasoning, has been detailed in Section 4.
Why Streaming Thinker Work?
Prior studies (fan2025missing; laban2025llms) have highlighted the risks of reasoning over incomplete inputs. This phenomenon reflects the inadequacy of batch-processing LLMs when facing local contexts, as their training objectives align with global visibility rather than incremental inference. However, Streaming Thinker circumvents these pitfalls through three distinct mechanisms. First, global information is deferred rather than lost. Unlike scenarios where critical conditions are permanently removed, our framework merely shifts the timing of acquisition, ensuring the model incorporates the full global context after the streaming phase. Second, we employ a conservative reasoning strategy. Instead of attempting premature complex reflection, the model concentrates on shallow reasoning” (e.g., intermediate calculations and entity tracking) during the streaming phase. This functions as an incremental pre-processing step that simplifies raw context. Third, this behavior is enforced via a specialized streaming training paradigm. Unlike batch models that rigidly apply full-context patterns to partial inputs—often falling into the over-thinking” trap—our model is explicitly adapted to local scopes, learning to process available information without jumping to erroneous conclusions.
| Metric | Batch Thinking Paradigm | Streaming Thinking Paradigm | ||||||
| Read | Prefill | Decoding | Read | Split KV | Prefill | Merge KV | Decoding | |
| Count (times) | 69.86 | 1 | 436.0 | 69.86 | 4.65 | 4.65 | 4.65 | 439.0 |
| Peak Mem (MB) | – | 7,991 | 7,998 | – | 7,957 | 7,993 | 7,940 | 7,999 |
| Total Mem (MB) | – | +222 | +102 | – | -289 | +215 | +217 | +103 |
| Bandwidth Cost (MB) | – | – | – | – | – | +29,173 | +434 | -3,310 |
| Avg Time (s) | 0.4 | 0.052 | 0.039 | 0.4 | 0.041 | 0.037 | ||
| Total Time (s) | 27.945 | 0.052 | 17.178 | 27.945 | 0.005 | 0.190 | 0.004 | 16.392 |
| First-token Latency (s) | 28.003 | – | – | 6.231 | – | – | – | – |
@@ 6 Related Work
Prior studies on efficient reasoning in LLMs have mainly focused on three directions: token compression (aggarwal2025l1; xia2025tokenskip; zhang2025lightthinker), structural quantization and pruning (liu2025quantization; srivastava2025towards; zhao2025skipgpt), and efficient decoding (pan2025learning; liao2025reward). Token compression methods such as Tokenskip (xia2025tokenskip) condense CoT into fewer tokens to reduce the inference cost. Structural approaches leverage quantization or pruning to compress model parameters for efficient reasoning such as (srivastava2025towards). Efficient decoding has been explored through parallel sampling of generation paths (pan2025learning) or by using smaller models for speculative prediction (liao2025reward) to reduce inference latency.
In contrast, our work introduces streaming processing as a complementary dimension of efficiency, allowing reasoning to proceed concurrently with input processing to reduce response latency.
@@
Recent research on streaming LLMs has focused on three directions: architecture adaptation (tong2025llm; raffel2024simultaneous; guo2024decoder), latency control (ma2018stacl; ahmed2025non; cheng2025seed), and modality adaptation (chen2024videollm; chen2025livecc; xie2024mini). Architecture adaptation addresses the mismatch between decoder-only transformers and streaming settings (tong2025llm), where issues such as positional interference and redundant re-encoding are mitigated through recurrent states, and modified attention mechanisms. Latency control emphasizes fine-grained alignment between input and output, ranging from fixed policies (e.g., wait-k (ma2018stacl)) to adaptive scheduling that optimize the accuracy–latency balance. Modality adaptation extends streaming LLMs beyond text to speech recognition (guo2024decoder), speech translation (cheng2025seed), and video understanding (chen2024videollm) by integrating modality-specific encoders and synchronization mechanisms. From a reasoning perspective, recent studies (xie2025interleaved; chiang2025stitch; xie2025mini) have explored alternating reasoning and generation to approximate streaming operation. Our work differs by explicitly modeling thinking while reading, enabling reasoning to evolve concurrently with incremental input.
@@ 7 Conclusion
Appendix A Motivation and Prospective Applications
A.1 Motivation and Prospective Applications
@@
The streaming thinking paradigm is also particularly well-suited for the challenges presented by embodied AI. An embodied agent, like a robot, operates within a dynamic physical environment, which distinguishes it from models that process static information. It continually receives multi-modal sensory data and is required to respond to this information in a timely manner. In this context, streaming thinking allows the agent to engage in continuous perceptual reasoning, interpreting incoming data to consistently update its internal model of the world. This capability supports dynamic instrumental reasoning, where the model can flexibly plan and re-plan its actions to navigate changing conditions and work towards a goal. For example, a household robot would need to reason about multiple factors at once, such as the movement of a person, the delicacy of an object it plans to handle, and the best path through a cluttered space, while adapting its motor commands accordingly. This close coupling of perception, reasoning, and action in a real-time loop is a core aspect of embodied cognition, and achieving it presents significant challenges for a conventional batch thinking approach.
Streaming Multimodal Understanding
Beyond text, streaming thinking is vital for interpreting continuous non-verbal data streams, such as live video feeds or audio environments. For instance, in video understanding, a model must reason about the temporal causality of events as they occur—identifying that an action in frame is a consequence of an event in frame . This is essential for applications like live sports commentary, real-time video surveillance for anomaly detection, or accessibility tools that provide live descriptions of the visual world for the visually impaired. By maintaining an evolving memory of the visual stream, the model can provide coherent, context-rich interpretations without the need to process the entire video offline.
A.2 Research Scope and Positioning
This work positions itself as a pioneering exploration into the paradigm of Streaming Thinking—technically enabling Large Language Models (LLMs) to perform concurrent reading and reasoning. Our primary contribution lies in establishing the architectural and training methodologies required to adapt LLMs for efficient, continuous data processing.
Regarding our evaluation scope, we utilize mathematical reasoning, logic reasoning, and contextual QA tasks primarily as a controlled testbed to verify the logical consistency of our streaming framework. These deterministic domains serve as a rigorous testbed to verify that our model can maintain logical consistency and accuracy under the strict constraints of streaming inference. Our future work aims to extend this paradigm to complex streaming scenarios.